Golden cup case study

You were provided with the following balance sheet for Golden Cup firm for the year ended Dec 31st, 2018.

Consolidated Balance sheet

Golden Cup.

As of Dec 31st, 2018

Assets

Liabilities + Owners Equity

Current Assets

Current Liabilities

Cash

40,000

Accounts Payable

12,000

Accounts Receivables

4,000

Notes Payable

6,000

Inventory

14,000

Accrue Wages

1000

Total Current Assets

58,000

Total Current Liabilities

19,000

Fixed Assets

Long term debt

40,000

Property, Plant, and Equipment

56,000

Owners’ equity

Goodwill

24,000

Common Shares

40,000

Total Fixed Assets

80,000

Retained Earnings

39,000

Total Owners equity

79,000

Total Assets

138,000

Liabilities + O.E

138,000

In addition to that, you know the following facts about firm’s operations throughout the year:

  • Golden Cup revenues for the year includes the following: Domestic revenues $160,000. International revenues $80,000. Out of Golden Cup’s sales, cost of sales and direct labor is 50% of annual revenues.

  • Because of the strong competition that it faces, Golden Cup has a generous marketing plan. Golden Cup signed a contract with the marketing planet Inc. by which the marketing agency will be responsible for Golden Cup marketing for five years period started this year. The contract costs Golden Cup $100,000 that were paid up front, however the company thinks this plan will affect its sales evenly over the five years period. Golden Cup also spends $30,000 in the form of general and administrative expenses per year. Golden Cup depreciable assets historical value is $40,000 and is depreciated on a straight line basis over 10 years.

  • Golden Cup pays interest rate of 10% on its Long-term debt outstanding.

  • Out of the year’s net income, Golden Cup is planning to repay $30,000 to its shareholders in the form of cash dividends. The company currently has 60,000 shares outstanding

Question 1

a- Please set up income statement for Golden Cup:

Consolidated Income Statement

Golden Cup.

As of Dec 31st, 2018

Show your workings here

Final answer here

Revenues

(-) Cost of goods sold

Gross margin

(-) Marketing expenses

(-) General and administrative expenses

(-) Depreciation

EBIT

(-) Interest expenses

EBT

(-) Tax expenses

Net income

Dividends

Additions to Retained Earnings
Answer #1

Consolidated Income Statement

Golden Cup.

As of Dec 31st, 2018

Show your workings here

Final answer here

Revenues

Domestic +International=160000+80000=240000 240000

(-) Cost of goods sold

=50% of 2400000=120000 -120000

Gross margin

120000

(-) Marketing expenses

=100000/5=20000 -20000

(-) General and administrative expenses

-30000

(-) Depreciation

-4000

EBIT

66000

(-) Interest expenses

-4000

EBT

62000

(-) Tax expenses

0*

Net income

62000

Dividends

30000

Additions to Retained Earnings

32000

*Tax details are not provided.